OpenStudy (anonymous):

if DR=43, what are the possible coordinates of point r if point d=-39?

OpenStudy (anonymous):

please help!

jimthompson5910 (jim_thompson5910):

I'm assuming you're on a number line. If D = -39, which point on this number line will be 43 units from D?

jimthompson5910 (jim_thompson5910):

oh sry, I mean "which points", there are two possible locations for R

OpenStudy (anonymous):

yea there are two and i said -43 but my teacher said it was wrong

OpenStudy (anonymous):

thanks you for helping me!

OpenStudy (anonymous):

-82*

jimthompson5910 (jim_thompson5910):

to find distance on a number line, say between two points x and y, you subtract and take the absolute value so |x-y| is the distance between x and y

jimthompson5910 (jim_thompson5910):

So the distance between D and R is |D - R|

OpenStudy (anonymous):

i know d = -39 but after that i get totally confused and dont know what i am doing

jimthompson5910 (jim_thompson5910):

this distance is given to be 43 units, so |D-R| = 43 and finally, D = -39, which means we can say |-39-R| = 43

jimthompson5910 (jim_thompson5910):

You just need to solve |-39-R| = 43 for R

OpenStudy (anonymous):

ok thanks :) you are awesome!

jimthompson5910 (jim_thompson5910):

np